# Geometric Analogues of Mie Scattering

### v1

@@ 3 Selected spatial geometries
in black, blue, green, orange, and red, respectively.
Once again, the limit $n\rightarrow\infty$ is somewhat difficult to process analytically, but from the previous graph, that limit is again clear.  As $n\rightarrow\infty$ the geometry is a cylinder of height $2H$ composed of 2-spheres each of radius $R$ (i.e. $\left[-H,H\right]\otimes S_{2}\left(R\right)$) attached at both ends to a disjoint pair of punctured $\mathbb{E}_{3}$s (i.e. $\mathbb{E}_{3}\smallsetminus\mathbb{B}_{3}\left(R\right)$ at $h=\pm H=\pm\alpha R$).
@@ 4 Amplitudes
Note the definition of the aspect ratio $y$ in terms of $H$ and $R$ for the wormhole differs from that for the foxhole by a factor of $2$.
@@ 8 Summary
Acknowledgements
TC received financial support from the United States Social Security Administration.
9 Appendix:  Perfect spheres
In black, individual $\sigma_{l}/\left(\pi R^{2}\right)$ for $l=0,1,2$ $3$ from left to right.
@@ 9 Appendix:  Perfect spheres
$\frac{1}{\pi R^{2}}\left(\sigma_{\text{TE perf cond sphere}}+\sigma_{\text{TM perf cond sphere}}\right)$ and
@@
$\frac{1}{\pi R^{2}}~{}\sigma_{\text{TE perf cond sphere}}$ and $\frac{2}{3}k^{4}R^{4}$ (light blue)
$\frac{1}{\pi R^{2}}~{}\sigma_{\text{TE perf cond sphere}}$ and $\frac{2}{3}k^{4}R^{4}$ (light blue)
@@
$l=1$ to $6$ summed contributions to $\sigma_{TE}$
$l=1$ to $6$ summed contributions to $\sigma_{TM}$.
@@
$l=1$ to $6$ summed contributions to $\sigma_{TE}+\sigma_{TM}$
@@
$l=1$ to $12$ summed contributions to $\sigma_{TE}+\sigma_{TM}$ in red and $\sigma_{P}$ in green
($l=1$ to $6$ summed contributions in orange).

### v2

@@ 3 Selected spatial geometries
[…]
@@ 4 Amplitudes
Note the definition of the aspect ratio $y$ in terms of $H$ and $R$ for the wormhole differs from that for the foxhole by a factor of 222.
@@ 6 Cross-sections
Figure 3: Fox Hole Scattering
Figure 4: Fox Hole Scattering
@@ 8 Summary
Acknowledgements TC received financial support from the United States Social Security Administration.
Disclosures The authors declare no conflicts of interest.
Appendix A Appendix:  Perfect spheres
Figure A-1: In black, individual $\sigma_{l}/\left(\pi R^{2}\right)$ for $l=0,1,2$ 333 from left to right.
@@ Appendix A Appendix:  Perfect spheres
Figure A-2: $\frac{1}{\pi R^{2}}\left(\sigma_{\text{TE perf cond sphere}}+\sigma_{\text{TM perf cond sphere}}\right)$ and
@@
Figure A-3: $\frac{1}{\pi R^{2}}~\sigma_{\text{TE perf cond sphere}}$ and $\frac{2}{3}k^{4}R^{4}$ (light blue)
Figure A-4: $\frac{1}{\pi R^{2}}~\sigma_{\text{TE perf cond sphere}}$ and $\frac{2}{3}k^{4}R^{4}$ (light blue)
@@
Figure A-5: $l=1$ to 666 summed contributions to $\sigma_{TE}$
Figure A-6: $l=1$ to 666 summed contributions to $\sigma_{TM}$.
@@
Figure A-7: $l=1$ to 666 summed contributions to $\sigma_{TE}+\sigma_{TM}$
@@
Figure A-8: $l=1$ to 121212 summed contributions to $\sigma_{TE}+\sigma_{TM}$ in red and $\sigma_{P}$ in green ($l=1$ to 666 summed contributions in orange).